# Can LLMs Reason Structurally? An Evaluation via the Lens of Data Structures

### v2

@@ 3.2 Prompt design
Example prompt for queue compound. A queue is a data structure in which items are added at one end and removed from the other, maintaining a first-in, first-out (FIFO) order. You should create a queue. There are two types of operations: (enqueue, k) adds k to the back. (dequeue) removes the front. You are given an empty queue initially. Q: What is the final queue after performing: (enqueue, 49) (dequeue) Answer the question in 8000 tokens.
For each task, we design a prompt template and populate it with synthetic data to produce individual problem instances. Each prompt follows a consistent format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final computed outcome. Following recommended practices in prompt engineering (OpenAI, 2025), we also append the instruction “Answer the question in <number> tokens” to guide models toward producing concise outputs within a specified token budget. We also implement five different prompting strategies for each task, as we detail in Section˜A.3.
3.3 Pipeline
@@
All data is synthetically generated to ensure efficiency, scalability, and minimal risk of training data contamination (Zhang et al., 2024a; Xu et al., 2024a). Each data structure and its operations are programmatically implemented to produce ground-truth outputs. Inputs are randomly sampled—numbers uniformly from 0–100 and strings from lowercase English letters. This modular approach enables accurate, reproducible evaluation and easy integration of new data structure types.
@@
The evaluation pipeline is fully objective and reproducible through automated and deterministic verification. This avoids subjective judging in proof-writing or math benchmarks (Chiang et al., 2024; Feuer et al., 2024; Ye et al., 2024). We integrate Structured Output using a predefined JSON Schema (natively supported by most models or via the Instructor library) and Pydantic, ensuring strict schema adherence. For example, a BST pre-order traversal must return a list[int]. Across 1,624 trials with nine models and six structures, we observed zero schema violations (Table˜33 in Appendix), confirming robustness of Structured Output.
@@
We compare model outputs to ground-truth answers, scoring 1 for correct and 0 for incorrect responses. Overall performance is reported as average accuracy across categories and difficulty levels from three runs. To ensure determinism, each task has a single well-defined solution. Any potential ambiguity (hash functions, tie-breaking rules) is explicitly specified in the prompts. We include an ablation with Levenshtein distance as a partial scoring metric in Section˜A.10.
4 Evaluation
@@
We evaluate ten state-of-the-art LLMs across two categories: instruction-tuned models and reasoning models. We select the latest flagship models from each major provider, with each problem evaluated three times (166,230 total evaluations). Instruction-tuned models are widely deployed due to their efficiency and scalability, making their structural reasoning capabilities practically important: we cover GPT-4.1-2025-04-14 (OpenAI et al., 2024), Llama3.3-70B (Grattafiori et al., 2024), Gemini-2.0-Flash-001 (Team et al., 2024), Claude-3-5-Sonnet-20241022 (Anthropic, 2024), and DeepSeek-V3 (DeepSeek-AI et al., 2025). Reasoning models are explicitly trained for complex, multi-step reasoning: we evaluate GPT-5-2025-08-07 with medium thinking effort (OpenAI, 2025), o4-mini-2025-04-16 (OpenAI, 2024), Gemini-2.5-Pro (stable) (Team et al., 2024), Claude-3-7-Sonnet-20250219 (Anthropic, 2024), and DeepSeek-R1 (DeepSeek-AI et al., 2025).
@@
We also study the impact of prompting strategies on data structure tasks. Unlike reasoning models with internal multi-step inference (e.g., reasoning tokens), instruction-tuned models are particularly sensitive to prompt formulation. We evaluate five prompting strategies: (i) Stepwise, which adds a “steps” field to the output JSON schema; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single example with intermediate reasoning steps; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text. See Section˜A.3 for examples.
@@ 4.1 Can LLMs understand and manipulate data structures?
In this section, we present our main experimental results on DSR-Bench, including main and its challenging subset challenge. We discuss insights from five instruction-tuned models in Section˜4.1.1 and from five reasoning models in Section˜4.1.2.
Table 1: Scores on DSR-Bench-main across ten LLMs. The table aggregates results by data structure and relationship type across three runs, including category scores and an overall score.
Relationship Data Structure GPT-5 (med) o4- mini Gemini- 2.5-Pro Claude-3. 7-Sonnet DeepSe ek-R1 GPT- 4.1 Gemini- 2.0-Flash Claude-3. 5-Sonnet DeepSe ek-V3 Llama 3.3-70B
Linear Array 1.00 1.00 1.00 0.99 0.99 0.94 0.90 0.96 0.98 0.69
Category avg. 1.00 1.00 1.00 0.99 0.99 0.94 0.90 0.96 0.98 0.69
Temporal Stack 1.00 1.00 1.00 0.97 0.99 0.55 0.36 0.99 0.41 0.04
Queue 1.00 1.00 1.00 1.00 0.98 0.55 0.36 0.99 0.43 0.25
LRU 1.00 1.00 1.00 0.30 0.16 0.85 0.78 0.82 0.01 0.50
Priority Queue 0.52 0.55 0.51 0.28 0.65 0.25 0.16 0.33 0.20 0.08
Category avg. 0.88 0.89 0.84 0.64 0.69 0.55 0.42 0.79 0.26 0.22
Associative Hashmap 0.87 0.51 0.28 0.63 0.33 0.06 0.10 0.16 0.00 0.00
Trie 0.94 0.68 0.62 0.08 0.49 0.18 0.17 0.49 0.02 0.00
Suffix Tree 0.98 0.73 0.90 0.91 0.96 0.00 0.01 0.08 0.67 0.00
Skip List 0.68 0.62 0.78 0.75 0.69 0.07 0.06 0.42 0.02 0.01
Category avg. 0.87 0.63 0.65 0.59 0.62 0.08 0.09 0.29 0.18 0.00
Hierarchical BST 1.00 0.86 0.97 0.64 0.73 0.59 0.43 0.71 0.58 0.34
Heap 0.61 0.68 0.38 0.40 0.48 0.20 0.10 0.27 0.15 0.07
RB tree 0.76 0.65 0.49 0.30 0.62 0.12 0.12 0.46 0.09 0.05
B+ tree 0.98 0.97 0.97 0.38 0.31 0.23 0.12 0.23 0.08 0.01
K-D Tree 0.85 0.47 0.59 0.00 0.45 0.00 0.01 0.03 0.00 0.00
K-D Heap 0.10 0.10 0.10 0.05 0.11 0.04 0.04 0.05 0.03 0.01
Category avg. 0.72 0.64 0.58 0.33 0.45 0.20 0.14 0.34 0.16 0.08
Network Graph 0.96 0.87 0.78 0.11 0.67 0.15 0.05 0.06 0.06 0.02
DSU 1.00 0.98 0.97 0.99 1.00 0.02 0.00 0.02 0.82 0.00
Geom Graph 0.16 0.16 0.13 0.02 0.13 0.07 0.02 0.02 0.01 0.02
Category avg. 0.71 0.67 0.63 0.38 0.60 0.08 0.03 0.03 0.30 0.01
Hybrid Bloom Filter 0.77 0.61 0.86 0.16 0.74 0.04 0.04 0.04 0.04 0.02
DAWG 0.34 0.25 0.23 0.06 0.14 0.05 0.06 0.07 0.06 0.01
Category avg. 0.56 0.43 0.55 0.11 0.44 0.05 0.05 0.06 0.05 0.02
Score Overall avg. 0.79 0.72 0.71 0.51 0.63 0.31 0.27 0.41 0.32 0.17
@@ 4.1.1 Instruction-tuned models
As shown in Table˜1, these models show sharp accuracy drops in tasks involving elements with multiple attributes. For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority. Similarly, in the hashmap task, manual inspection of errors shows that models confuse keys and values, delete the wrong items, or hallucinate entries. These results reveal a key limitation for real-world deployment, where managing entities with multiple interacting properties, such as deadlines, priorities, or key-value records, is common.
@@
We see from Table˜1 that, while models perform reliably on BST, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across ancestral levels. Performance declines further on B+ trees, which requires reasoning over wider spans of child pointers, and on graph traversal tasks with many-to-many relationships. Manual inspection of reasoning traces reveals failures to retain earlier information: in Geometric Graph-long, all GPT-4.1 errors stemmed from dropping nodes during intermediate steps.
Table 2: Average scores across all tasks for instruction-tuned models under different prompting strategies.
Model Stepwise 0-CoT CoT 3-shot None
DeepSeek-V3 0.67 0.67 0.66 0.64 0.55
Llama-3.3 0.46 0.46 0.48 0.46 0.34
GPT-4.1 0.82 0.83 0.94 0.80 0.59
Gemini-2.0-Flash 0.57 0.86 0.59 0.87 0.58
Claude-3.5-Sonnet 0.72 0.92 0.72 0.93 0.69
@@
As shown in Table˜2, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning can be beneficial. Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Section˜A.5); (ii) Structured prompts like CoT and 3-shot are most effective for uncommon data structures, but require careful design. A representative case is suffix-tree: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Section˜A.5). We include more analysis of CoT to provide meaningful takeaways for practitioners in Section˜A.6.
@@ 4.1.2 Reasoning models
Table 3: Scores on the challenge suite for five reasoning models.
Model Score Priority Queue Suffix Tree Trie Skip List Heap Red-bla ck tree B+ Tree K-D Tree K-D Heap DSU Geom Graph Bloom Filter DAWG
GPT-5 0.50 0.52 0.98 0.92 0.51 0.71 0.585 0.98 0.85 0.10 1.00 0.19 0.47 0.00
Gemini-2.5-Pro 0.47 0.24 0.89 0.59 0.61 0.62 0.60 0.96 0.67 0.00 0.99 0.00 0.69 0.00
DeepSeek-R1 0.36 0.48 0.90 0.05 0.54 0.27 0.37 0.21 0.01 0.01 0.99 0.01 0.31 0.00
o4-mini 0.34 0.30 0.37 0.32 0.41 0.71 0.37 0.94 0.38 0.00 0.94 0.00 0.07 0.06
Claude-3.7-Sonnet 0.21 0.04 0.79 0.00 0.61 0.13 0.12 0.13 0.00 0.00 0.98 0.00 0.00 0.00
@@
From Table˜1, we see reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures. However, the overall score remains below 0.5 on challenge in Table˜3, in particular for long tasks and complex data structures. For example, the highest score on skip list is only 0.61, despite its prevalence in introductory-level textbooks and its wide use in dictionaries and maps. Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Section˜A.4). To further probe these limitations, we introduce spatial, described in Section˜4.2.
@@
In an ablation on K-D Heap in Table˜5, switching the tie-breaking rule from lexicographic order to Euclidean norm causes o4-mini’s score to drop by over 0.40, as it continues to apply lexicographic comparisons. Directly querying o4-mini about the default implementation of a K-D heap confirms this point: it assumes lexicographic keys. These results suggest that reasoning models such as o4-mini may struggle to override entrenched priors learned from training, limiting their reliability on tasks with user-defined constraints.
@@ 4.2 Can LLMs reason structurally on spatial data?
Real-world data is often represented in high-dimensional feature spaces. To assess whether LLMs can reason over such spatial data, we extend the benchmark with the spatial suite, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space. These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics. Given the complexity of these tasks, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps.
Table 4: Scores for the three spatial data structures with input data of varying dimensionality ($k=1,2,3,5$).
K-D Heap K-D Tree Geom. Graph
$k$ GPT-4.1 o4-mini GPT-4.1 o4-mini GPT-4.1 o4-mini
1 $0.74$ $0.82$ $0.91$ $0.82$ $0.18$ $0.88$
2 $0.30$ $0.34$ $0.86$ $0.69$ $0.04$ $0.90$
3 $0.26$ $0.26$ $0.92$ $0.68$ $0.01$ $0.76$
5 $0.21$ $0.21$ $0.73$ $0.64$ $0.00$ $0.71$
Table 5: Performance on the two comparison metrics for K-D Heap with different $k$ values.
$k$ Lexicographic Euclidean
1 0.79 0.82
2 0.87 0.34
3 0.84 0.26
5 0.90 0.21
Performance degrades as input dimensionality increases.
As shown in Table˜5, accuracy declines for both models as dimensionality increases. Higher-dimensional data challenges models with more complex computation over distance metrics and partitions, limiting their effectiveness in spatial tasks. For instance, K-D trees are widely used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision (Silpa-Anan & Hartley, 2008). Interestingly, 2D outperforms 1D in Geometric Graph, likely due to its more common presence in training data from textbooks.
Table 6: Scores on K-D Tree with varying input data distributions.
Distribution GPT-4.1 o4-mini
Uniform $0.86$ $0.69$
Moon $0.42$ $0.62$
Blob $0.33$ $0.62$
Circle $0.31$ $0.67$
@@
We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data. We test K-D tree construction tasks using three non-uniform distributions from scikit-learn (Pedregosa et al., 2011): circles, moons, and blobs (illustration in Figure˜3, Section˜A.7). As shown in Table˜6, GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in the training data. Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning. In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts. A more in-depth inspection of errors and discussion on root causes can be found in Section˜A.7.
4.3 Can LLMs reason structurally on natural language tasks?

### v3

@@ 3.3 Prompt design
Example prompt for queue compound.
(i) A queue maintains a first-in, first-out (FIFO) order, where items are added at one end and removed from the other.
(ii) There are two operations: (enqueue, k) adds $k$ to the back; (dequeue) removes the front.
(iii) You are given an empty queue initially.
(iv) Q: What is the final queue after: (enqueue, 49), (enqueue, 85), (dequeue), …
For each task, we design a prompt template and populate it with synthetic data to generate problem instances. Each prompt has the following format: (i) a concise description of the data structure; (ii) a detailed explanation of the operations to be performed, written to avoid ambiguity; (iii) the initial state of the data structure (e.g., an existing tree or list) and any additional inputs required to execute the task (e.g., new elements to insert or delete); and (iv) a specific question requesting the final outcome. Following recommended practices in prompt engineering, we also append the instruction “Answer the question in <number> tokens” to encourage concise outputs within a specified token budget. We also implement five different prompting strategies, as we detail in the next section.
3.4 Data generation and evaluation
@@
As a key strength of DSR-Bench, all data is synthetically generated to ensure scalability and minimal risk of data contamination (Zhang et al., 2024a; Xu et al., 2024a). Numerical inputs are sampled uniformly from the range of 0 to 100, and string inputs are composed of uniformly sampled lowercase English letters. Each data structure and its operations are programmatically implemented to produce ground-truth outputs, enabling accurate labeling and easy integration of new data structures.
@@
The automated evaluation is fully deterministic and reproducible, constituting another notable advantage of DSR-Bench. This feature avoids subjective human or LLM-based judging, as commonly required in proof-writing or math benchmarks (Chiang et al., 2024; Feuer et al., 2025; Ye et al., 2025). We use Structured Output (either built-in for most models or supported via Ollama/Instructor) to enforce outputs that conform to a predefined JSON Schema. For example, a BST pre-order traversal must return a list[int]. Across 1,624 trials with nine models and six structures, we observe zero schema violations (Table˜30 in the Appendix), confirming its robustness.
@@
We use binary (0/1) scoring, comparing against a single well-defined solution. Any potential ambiguity (hash functions, tie-breaking rules) is explicitly specified in the prompts. As an alternative metric, we include an ablation using the Levenshtein distance as a partial-credit metric in Appendix˜I, which shows that binary scoring provides a clearer distinction and a fairer comparison.
4 Empirical analysis
@@
We evaluate 13 state-of-the-art LLMs, including instruction-tuned and reasoning models, as well as open- and closed-source models. We select the flagship models from each major provider, with each problem evaluated three times (166,230 total evaluations). Instruction-tuned models are widely deployed due to their efficiency and scalability, making their structural reasoning capabilities practically important: we cover GPT-4.1-2025-04-14 (OpenAI, 2025a), Gemini-2.0-Flash-001 (Google, 2025), DeepSeek-V3 (DeepSeek-AI et al., 2025), Qwen3-8B-Instruct (Yang et al., 2025), Claude-3-5-Sonnet-20241022 (Anthropic, 2024), Llama3.3-70B (Grattafiori et al., 2024), and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Reasoning models are explicitly trained for complex, multi-step reasoning: we evaluate GPT-5-2025-08-07 with medium thinking effort (OpenAI, 2025b), o4-mini-2025-04-16 (OpenAI, 2025c), Gemini-2.5-Pro (stable) (Comanici et al., 2025), Claude-3-7-Sonnet-20250219 (Anthropic, 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), Phi-4-reasoning-14B (Abdin et al., 2025).
@@
We also study the impact of prompting strategies on data structure tasks. Unlike reasoning models with internal multi-step inference (e.g., reasoning tokens), instruction-tuned models are particularly sensitive to prompt formulation. We evaluate five prompting strategies: (i) Stepwise, which adds a “steps” field to the output JSON schema; (ii) 0-CoT, which appends “Let’s think step by step” without examples; (iii) CoT, which provides a single example with intermediate reasoning steps; (iv) 3-shot, which includes three input-output examples; and (v) None, the default prompting setting with no added text. See Appendix˜C for examples.
@@ 4.1 Can LLMs understand and manipulate data structures?
We present empirical results on DSR-Bench-main and its challenging subset challenge. We discuss insights from seven instruction-tuned models in Section˜4.1.1 and from six reasoning models in Section˜4.1.2.
Table 2: Scores on DSR-Bench-main across 13 models. The table aggregates results by data structure and relationship type across three runs, including category scores and an overall score.
Relationship Data Structure GPT-5 (med) Gemini- 2.5-Pro o4- mini DeepSe ek-R1 Claude-3. 7-Sonnet Claude-3. 5-Sonnet Qwen3-8 B-Instruct DeepSe ek-V3 GPT- 4.1 Gemini- 2.0-Flash Llama 3.3-70B Phi-4- R-14B Mixtral- 8x7B
Linear Array 1.00 1.00 1.00 0.99 0.99 0.96 0.99 0.98 0.94 0.90 0.69 0.63 0.49
Category avg. 1.00 1.00 1.00 0.99 0.99 0.96 0.99 0.98 0.94 0.90 0.69 0.63 0.49
Temporal Stack 1.00 1.00 1.00 0.99 0.97 0.99 0.99 0.41 0.55 0.36 0.04 0.12 0.09
Queue 1.00 1.00 1.00 0.98 1.00 0.99 0.97 0.43 0.55 0.36 0.25 0.05 0.03
LRU 1.00 1.00 1.00 0.16 0.30 0.82 0.49 0.01 0.85 0.78 0.50 0.00 0.00
Priority Queue 0.52 0.51 0.55 0.65 0.28 0.33 0.35 0.20 0.25 0.16 0.08 0.01 0.01
Category avg. 0.88 0.84 0.89 0.69 0.64 0.79 0.70 0.26 0.55 0.42 0.22 0.04 0.03
Associative Hashmap 0.87 0.28 0.51 0.33 0.63 0.16 0.01 0.00 0.06 0.10 0.00 0.00 0.02
Trie 0.94 0.62 0.68 0.49 0.08 0.49 0.00 0.02 0.18 0.17 0.00 0.00 0.00
Suffix Tree 0.98 0.90 0.73 0.96 0.91 0.08 0.36 0.67 0.00 0.01 0.00 0.11 0.00
Skip List 0.68 0.78 0.62 0.69 0.75 0.42 0.03 0.02 0.07 0.06 0.01 0.01 0.00
Category avg. 0.87 0.65 0.63 0.62 0.59 0.29 0.10 0.18 0.08 0.09 0.00 0.03 0.00
Hierarchical BST 1.00 0.97 0.86 0.73 0.64 0.71 0.52 0.58 0.59 0.43 0.34 0.18 0.09
Heap 0.61 0.38 0.68 0.48 0.40 0.27 0.29 0.15 0.20 0.10 0.07 0.03 0.01
RB tree 0.76 0.49 0.65 0.62 0.30 0.46 0.09 0.09 0.12 0.12 0.05 0.02 0.01
B+ tree 0.98 0.97 0.97 0.31 0.38 0.23 0.14 0.08 0.23 0.12 0.01 0.03 0.00
K-D Tree 0.85 0.59 0.47 0.45 0.00 0.03 0.00 0.00 0.00 0.01 0.00 0.00 0.00
K-D Heap 0.10 0.10 0.10 0.11 0.05 0.05 0.05 0.03 0.04 0.04 0.01 0.01 0.01
Category avg. 0.72 0.58 0.64 0.45 0.33 0.34 0.18 0.16 0.20 0.14 0.08 0.05 0.02
Network Graph 0.96 0.78 0.87 0.67 0.11 0.06 0.13 0.06 0.15 0.05 0.02 0.01 0.01
DSU 1.00 0.97 0.98 1.00 0.99 0.02 0.73 0.82 0.02 0.00 0.00 0.16 0.02
Geom Graph 0.16 0.13 0.16 0.13 0.02 0.02 0.00 0.01 0.07 0.02 0.02 0.00 0.00
Category avg. 0.71 0.63 0.67 0.60 0.38 0.03 0.29 0.30 0.08 0.03 0.01 0.06 0.01
Hybrid Bloom Filter 0.77 0.86 0.61 0.74 0.16 0.04 0.12 0.04 0.04 0.04 0.02 0.01 0.00
DAWG 0.34 0.23 0.25 0.14 0.06 0.07 0.00 0.06 0.05 0.06 0.01 0.00 0.00
Category avg. 0.56 0.55 0.43 0.44 0.11 0.06 0.06 0.05 0.05 0.05 0.02 0.00 0.00
Score Overall avg. 0.79 0.71 0.71 0.63 0.51 0.41 0.39 0.32 0.31 0.27 0.17 0.13 0.09
@@ 4.1.1 Instruction-tuned models
Figure 3: Average scores across all tasks for five instruction-tuned models under different prompting strategies.
@@
As shown in Table˜2, these models show sharp degradation in tasks involving elements with multiple attributes. For instance, while they perform well on queue, their accuracy drops 30-50% on priority queue, where each element includes a priority. Similarly, in the hashmap task, manual inspection of errors shows that models confuse keys and values, delete the wrong items, or hallucinate entries. These results reveal a key limitation for real-world deployment, where managing entities with multiple properties, such as deadlines and key-value records, is common.
@@
We see from Table˜2 that, while models perform reliably on BST, their accuracy drops by over 30% on red-black trees, reflecting the difficulty of handling multi-hop properties such as maintaining balance across ancestral levels. Performance declines further on B+ trees, which requires reasoning over wider spans of child pointers, and on graph traversal tasks with many-to-many relationships. Manual inspection of reasoning traces reveals failures to retain earlier information: in Geometric Graph-long, for example, all GPT-4.1 errors stemmed from dropping nodes during intermediate steps.
@@
As shown in Figure˜3, the None prompt performs worst, suggesting that prompts encouraging stepwise reasoning is generally beneficial. Our findings indicate two practical strategies: (i) Lightweight prompts such as Stepwise and 0-CoT are easily implemented and consistently improve performance (Appendix˜E); (ii) Crafted prompts like CoT and 3-shot are most effective for uncommon data structures, but require careful design. A representative case is suffix-tree: across all models, zero-shot accuracy is below 0.40, but a well-designed CoT prompt doubles accuracy for three models (Appendix˜E). We include additional CoT analysis with practical takeaways in Section˜E.1.
@@ 4.1.2 Reasoning models
Table 3: Scores of the challenge subset for reasoning models.
Task GPT-5 o4-mini Gemini- 2.5-Pro DeepSeek -R1 Claude-3.7 -Sonnet
Challenge Score 0.46 0.32 0.30 0.21 0.10
Priority Queue 0.28 0.30 0.23 0.48 0.04
Skip List 0.51 0.41 0.53 0.54 0.61
Heap 0.71 0.71 0.53 0.27 0.13
Red-black tree 0.59 0.37 0.08 0.37 0.12
B+ Tree 0.98 0.94 0.94 0.21 0.13
K-D Tree 0.67 0.38 0.16 0.01 0.00
K-D Heap 0.00 0.00 0.00 0.01 0.00
Geom Graph 0.19 0.00 0.02 0.01 0.00
Bloom Filter 0.47 0.07 0.69 0.31 0.00
DAWG 0.00 0.06 0.01 0.00 0.00
Hashmap 0.71 0.26 0.11 0.12 0.04
@@
From Table˜2, we see reasoning models outperform instruction-tuned models in general, especially on hierarchical and networked structures. However, the overall score remains below 0.5 on challenge in Table˜3, in particular for long tasks and complex data structures. For example, the highest score on skip list is only 0.61 for the challenge subset, despite its prevalence in introductory-level textbooks and its wide use in dictionaries and maps. Notably, accuracy on K-D Tree, K-D Heap, and Geometric Graphs is low even for short tasks, suggesting that high-dimensional spatial reasoning remains a significant challenge (Appendix˜D). To further probe these limitations, we introduce spatial, described in Section˜4.2.
@@
Ablation on K-D Heap (Figure˜5(a)) shows that switching the tie-breaking rule from lexicographic order to Euclidean norm leads to a drop of over 0.40 on o4-mini. We observe that the model continues to apply lexicographic order, even when instructed otherwise. When queried directly, o4-mini confirms that it assumes lexicographic order for K-D heaps by default. These results suggest that o4-mini struggles to adapt to user-defined constraints, likely due to its reliance on learned priors from training.
@@ 4.2 Can LLMs reason structurally on spatial data?
Real-world data is often represented in high-dimensional feature spaces. To assess whether LLMs can reason over such spatial data, we extend the benchmark with the spatial probe, which includes three multi-dimensional variants: K-D Heap, K-D Tree, and Geometric Graph embedded in Euclidean space. These structures are common in practice; for instance, K-D trees are key data structures in computer vision and graphics. Given the complexity of these tasks, we use GPT-4.1 with the Stepwise prompt to encourage intermediate reasoning steps.
Figure 4: Scores for the three spatial data structures with input data of varying dimensionality ($k=1,2,3,5$).
Performance degrades as dimensionality increases.
In Figure˜4, accuracy declines for both models as dimensionality increases. Higher-dimensional data challenges models with complex computation over distance metrics and partitions, limiting their effectiveness in spatial tasks. For instance, K-D trees are widely used to expedite nearest neighbor queries over 128-dimensional SIFT descriptors in computer vision (Silpa-Anan and Hartley, 2008). Interestingly, 2D outperforms 1D in Geometric Graph, likely because it is more common in training data from textbooks.
Figure 5: (a) Performance on the two metrics for K-D Heap. (b) K-D Tree with varying input data distributions.
@@
We assess LLM robustness to distribution shifts by comparing performance on uniformly sampled versus skewed or clustered data. We test K-D tree construction tasks using three non-uniform distributions from scikit-learn (Pedregosa et al., 2011): circles, moons, and blobs (illustration in Figure˜8, Appendix˜F). As shown in Figure˜5(b), GPT-4.1’s performance drops sharply on non-uniform inputs, possibly due to a higher likelihood of uniformly distributed examples in the training data. Since task difficulty is held constant, this gap suggests a reliance on pattern memorization rather than true reasoning. In contrast, o4-mini shows a smaller drop, indicating that reasoning models may generalize better to distribution shifts. A more in-depth inspection of errors and discussion on root causes can be found in Appendix˜F.
4.3 Can LLMs reason structurally on realistic tasks?